The angular momentum of a flywheel having a Moment of inertia 0.125kgm^3 decreases from 3.0 to 2.0kgm^2/s^2 in a period of 1.5s
(i) Assuming a uniform angular acceleration, through how many revolutions will the flywheel have tuned?
(ii) how much work is done?
1
Expert's answer
2020-06-25T09:35:28-0400

Solution.

"I=0.125kgm^2;"

"L1=3.0kgm^2\/s;"

"L_2=2.0kgm^2\/s;"

"t=1.5s;"

i)"M=\dfrac{\Delta L}{\Delta t};"

"M=\dfrac{1kgm^2/s}{1.5s}=0.7kgm^2/s;"

"M=I\epsilon\implies \epsilon=\dfrac{M}{I};"

"\epsilon=\dfrac{0.7kgm^2/s}{0.125kgm^2}=5.6rad/s^2;"

"\epsilon=\dfrac{\Delta\omega}{\Delta t};" "\Delta \omega=\dfrac{\Delta\phi}{\Delta t};"

"\Delta\phi=\epsilon(\Delta t)^2;"

"\Delta\phi=5.6rad/s^2\sdot(1.5s)^2=12.6rad;"

"n=\dfrac{12.6rad}{2\pi rad}=8;"


ii )"A=\dfrac{I\omega_1^2}{2}-\dfrac{Iw_2^2}{2};"

"L_1=I\omega_1;"

"\omega_1=\dfrac{3.0kgm^2/s}{0.125kgm^2}=24rad/s;"

"\omega_2=\dfrac{2.0kgm^2/s}{0.125kgm^2}=16rad/s;"

"A=\dfrac{0.125kgm^2\sdot(24rad/s)^2}{2}-\dfrac{0.125kgm^2\sdot(16rad/s)^2}{2}=20J;"

Answer: "i)n=8;"

"ii)A=20J."